OBSERVING A COLUMN-DEPENDENT ζ IN THE HORSEHEAD PDR

\underline{P.B. RIMMER}, Department of Physics, Ohio State University, Columbus, OH 43210; O. MORATA, Institute of Astronomy and Astrophysics, Academia Sinica, Tapei 11677, Taiwan; E. ROUEFF, Observatoire de Paris, LUTH and Université Denis Diderot, Place J. Janssen 92190 Meudon, France; E. HERBST, Departments of Astronomy, Physics, and Chemistry, Ohio State University, Columbus, OH 43210.

The molecules CCH, C_{4}H, HCO^{+}, and HC_{3}N have been observed at the edge of the Horsehead Nebula in abundances far higher than those predicted by Photodissociation Region (PDR) models^{a}. Using a column-dependent cosmic ray ionization rate (ζ)^{b}, we model the edge of the Horsehead Nebula as a one-dimensional nearly-“edge-on” heterogeneous PDR with temperature ranging from 15-250 K, number density from 10^{3} - 5 × 10^{5} cm^{−3}, and ζ = 10^{−16} - 10^{−14} s^{−1}. The resulting abundances for the molecules listed above are much closer to the observed abundances. In this talk, we will discuss this method, its results, and the usefulness of incorporating a column-dependent ζ in astrochemical PDR models, especially in the advent of ALMA.

aJ. Pety et al., A&A \underline{435}, 885 (2005)

bM. Padovani, D. Galli, A.E. Glassgold, A&A \underline{501}, 619 (2009)